Assume that a car decelerates at 2.0m/s2 and comes to a stop after traveling 25m.

a) Find the speed of the car at the start of the deceleration


1
Expert's answer
2022-03-06T18:05:40-0500

The speed of the car at the start of the deceleration

"v_i=\sqrt{2ad}\\
=\rm\sqrt{2*2.0\: m/s^2*25\:m}=10\: m/s"